# Robot-mediated physical Human-Human Interaction in Neurorehabilitation: a position paper

### v1

@@ IV RHHI Potential in NR
RHHI combines the advantages of robotics (Sec. IV-A) while incorporating therapist expertise into the control (Sec. IV-B). This section summarizes both contributions (Tab. I).
@@ IV-A Added Value of Robotics
Robotic devices complement therapists by partially offloading the patient’s weight and accurately transmitting the therapist’s intended movements to the patient [19]. This feature is particularly valuable during prolonged and intensive rehabilitation sessions that require continuous physical effort. It is especially beneficial for patients with severe motor impairments, as therapists must maintain patient posture, joint alignment, and smooth movements throughout an exercise. Moreover, the high repeatability of robotic devices can relieve therapist effort by replicating movements from previous demonstrations.
@@
In addition, robotic platforms enrich rehabilitation data by capturing simultaneous kinematic and dynamic information from both the patient and the therapist [50]. Kinematic data provide movement trajectories, joint angles, and velocities, revealing patterns and achievable ranges of motion. Dynamic data, such as interaction forces, account for the level of assistance or resistance provided and patients’ kinetic responses, giving deeper insight into the quality of the interaction.
@@
RHHI holds great potential beyond clinical settings. In recent years, the availability of commercially-available assistive devices intended for domestic use has been increasing [51]. Applying RHHI to these devices could enable continued support for exercises outside of the clinic. For individuals requiring continuous care, RHHI enables remote assistance and rehabilitation, allowing caregivers and therapists to monitor activities, provide real-time support, and potentially ensure a more effective continuum of care.
IV-B Added Value of Therapists
@@
In contrast to traditional robotic systems, which typically rely on predefined trajectories and often require time-consuming calibration, RHHI allows exercise customization and adjustment on-the-fly by incorporating therapist expertise. Moreover, traditional systems often require settings to be adjusted by non-technical users, like clinicians, who may not fully understand their impact, potentially limiting the effectiveness and usability of these systems. Instead, in RHHI, robotic trajectories are dynamically shaped by the real-time input of another user, such as a physical therapist, enhancing adaptability and intuition in robotic interventions [52, 53]. During each session, therapists can flexibly select and modify exercises or movements based on immediate and precise patient assessments without having to set predefined trajectories.
@@
Therapists can also modulate assistance levels throughout rehabilitation. They may initially offer substantial support, provide transparent supervision during assessments, or introduce resistance to encourage active engagement. Although quantifying assistance or resistance can be difficult, therapists can intuitively assess these needs through direct interaction and visual observation, as in conventional therapy [54, 55].
@@ V RHHI Applications in NR
The potential applications of RHHI in NR are diverse and far-reaching. In this section, we outline several practical use cases, starting with conventional manual therapy examples and extending to advanced robot-mediated approaches. The section begins by presenting a potential implementation of RHHI for impairment assessment (Sec. V-A), followed by three different training examples (Sec. V-B, V-C, V-D). Finally, we explore the use of RHHI in group therapy (Sec. V-E) and remote therapy (Sec. V-F).
@@ V-A Motor Assessment
Assessing patient impairment levels is crucial in rehabilitation [50, 56]. A thorough evaluation must capture four domains: pathology, impairment, participation, and contextual factors [57]. When such information is recorded in a standardized, well-structured format, this enriches insight into the rehabilitation process, streamlines data sharing and reuse, and ultimately delivers more cost- and time-efficient care [58]. Assessing physiological changes (Box 1) in the brain or spinal cord to understand motor recovery can be time-consuming and is often not feasible in many clinics. Therefore, higher-level impairments such as muscle weakness, spasticity, and motor coordination are often evaluated and used as markers for recovery. Functional task scales, such as the 9-hole peg test and Box & Blocks, offer valuable information about a person’s ability to perform daily activities [59]. Because these metrics can change over time, continuous monitoring is crucial to adapt therapy to patient needs.
@@
Combining the quantitative precision of robotic devices with the personalized insights of manual assessments through RHHI enables a comprehensive and adaptive evaluation approach. An innovative RHHI example is the “Beam-Me-In” strategy presented by Baur et al. [54]. Here, a therapist and patient are connected through exoskeleton robots, enabling therapists to directly perceive patient limitations (e.g., muscle tone, range of motion). This embodied experience provides therapists with a more accurate understanding of patient impairments compared to traditional observation alone. RHHI simultaneously quantifies joint kinematics and interaction forces, enhancing the precision of impairment assessments and treatment adjustments. At the same time, the therapist can design assessment movements in real time, informed by clinically relevant parameters—such as perceived effort, fatigue, or pain—that are difficult to quantify objectively.
@@
Box 1: Physiological Signal and RHHI Electromyography (EMG) plays a crucial role in neurorehabilitation by quantifying volitional input during RHHI. EMG has been used to study muscle co-contraction mechanisms [60, 24], where antagonistic muscles activate simultaneously to stabilize joints during haptic interactions. Analyzing EMG signals helps researchers understand how individuals modulate muscle activity in response to a partner’s movements and task-related uncertainties [31]. In teleimpedance control [61], EMG provides real-time muscle activity data, allowing robots to dynamically adjust impedance. Future research should explore integrating teleimpedance into RHHI to enable bidirectional control of both joint posture and user impedance. In combination with Functional Electrical Stimulation (FES) can ensure synchronized movement between therapist and patient. For instance, therapist EMG signals could be used to trigger FES stimulation on the patient, reinforcing the relearning of movement patterns [62]. Electroencephalography (EEG) can be used to assess and quantify Interpersonal Synchronization, and simultaneous EEG measurements, known as EEG hyperscanning, have been used to provide insights into neural synchronization between individuals engaged in shared activities [63]. Research has shown that inter-brain synchronization, particularly in regions associated with the mirror neuron system [64], increases during coordinated activities such as gesture imitation [65] and finger movements [66]. Autonomic Nervous System response, such as respiration rate and skin conductance, offers alternative markers for interpersonal engagement. These signals have been used to assess task difficulty and enjoyment during interactive activities [67, 68] and robot-assisted therapy [69, 70, 71, 72]. In RHHI, these techniques can provide indicators of motivation and engagement.
@@ V-B Upper-Limb Training
Upper-limb rehabilitation is essential for individuals undergoing motor recovery, as impairments in motor control and weakness of the hand, wrist, elbow, and shoulder are common. Several studies have investigated RHHI during upper-limb training, showing that healthy individuals track movement trajectories more accurately when haptically connected with a partner compared to training alone. These improvements have been linked to estimation of the partner’s movement strategy to refine their own movement control [73, 24, 34].
Despite promising findings, few existing RHHI studies have evaluated these approaches in individuals with neurological impairments. Some studies have involved patient populations performing social rehabilitation games [74, 38], however, the infrastructure used in these studies did not allow users to physically interact. Both studies emphasized social interaction to boost engagement during training, highlighting the importance of the interaction scenario in future RHHI work. The only notable exception is the study by Waters et al. [75]. This work presented preliminary results on three chronic stroke participants paired with a healthy partners during a trajectory tracking task, showing improved individual motor learning as a result of haptically interacting with an unimpaired peer and thus validating the feasibility of such an approach for upper-limb training. Future work could extend these RHHI findings during trajectory tracking tasks to larger cohorts in longitudinal studies, to determine the effects of such training on sustained functional benefits.
One promising research direction has been suggested by early studies on human–human interaction during handwriting [76, 77]. To date, RHHI research has primarily focused on arm-level tasks, with limited attention to hand-specific activities requiring dexterous manipulation. Although various hand rehabilitation robots have been developed [78, 79, 80], none currently support direct haptic interaction between users. Enabling such interaction could significantly expand the range of functional and dexterity-based tasks that patients can perform in collaboration with therapists.
@@ V-C Sit-to-Stand Training
The ability to transition from sitting to standing (StS) is fundamental for independent living. It is one of the most common functional tasks and correlates with clinical balance scores in stroke patients [81]. Post-stroke individuals often struggle with StS due to deficits in strength, motor control, and proprioception, leading to asymmetric force distribution, slower movements, and greater pressure displacement [15].
Vianello et al. [82] introduced a dual-exoskeleton system that allows a therapist and patient to interact physically across multiple joints and contact points. This setup enables the therapist to guide the patient while incorporating programmable support features, such as balance assistance and safety measures essential for StS tasks. The system was tested with healthy individuals performing synchronized StS movements using shared visual feedback. To reduce fatigue during prolonged sessions, the exoskeletons offload part of the patient’s body weight, allowing them to focus on coordination rather than physical exertion. The study also explored two interaction configurations (Sec. III-A): (1) joint-level coupling, where a spring-damper system links the joint configurations of both users, and (2) task-space coupling, where a virtual connection between their centers of mass allows for flexible posture while preserving overall coordination.
By combining robotic assistance with therapist guidance, RHHI could provide a more adaptive and effective approach to StS training, improving patient outcomes while reducing therapist strain. This approach could be beneficially extended to standing balance training using a similar setup.
@@ V-D Gait Training
Conventional gait therapy typically involves therapists manually guiding patients to improve balance and gait patterns, often using partial body-weight support. During treadmill-based training, the therapist usually sits in front of the patient and manually assists leg movements. In overground training, body-weight support is also commonly employed, and multiple therapists may be required to support and guide both legs [83]. Conventional gait therapy may include tasks such as stair climbing and obstacle avoidance; however, these require specialized body-weight support systems to safely assist patients during more complex maneuvers: equipment that is not available in all clinical settings. Robot-Assisted Gait Training (RAGT) supports patients during high-intensity training [19, 84], but it typically facilitates only a limited range of exercises—most often treadmill walking, and less frequently, overground walking.
@@
RHHI offers adaptive, interactive exercises tailored to individual patient needs and the possibility to extend RAGT to a wider range of exercises without prior trajectory definition. Amato et al. [53] proposed and validated a system in which a therapist wearing an Inertial Measurement Unit (IMU)-based suit provides unidirectional haptic feedback to a patient in an exoskeleton during treadmill walking and obstacle avoidance. Koh et al. [85] virtually connected a robotic manipulandum to a robotic arm physically attached to the paretic leg of post-stroke patients. This setup enables bidirectional interaction, allowing therapists to actively influence the patient’s leg movements and demonstrating improved kinematics during training.
@@
Both approaches [53, 85] support collaborative interaction and assist patients in forming internal movement models through therapist-guided feedback. While only Amato et al. demonstrated obstacle avoidance, both systems have the potential to support a wider range of exercises. The main limitation lies in hardware constraints; for example, Koh et al. use a grounded manipulator to assist the paretic leg, which may restrict its use in overground activities such as walking, stairs, or obstacles. Future developments could explore mobile manipulators to extend support beyond treadmill-based training (see Sec. VI for future technical challenges). A key distinction between the two examples proposed is that the method by Amato et al. uses a unidirectional connection—only the patient perceives haptic feedback—which lowers cost and setup time [53] (see Sec.VII). Another difference is that the approach of Amato et al. supports interaction with both patient limbs, while Koh et al. focuses on a single limb. Koh et al. tested their system with hemiparetic stroke patients, targeting the impaired limb, but adapting the method for other populations, such as those with SCI, would require further development [85].
@@ V-E Group Therapy
Group therapy is widely used in rehabilitation, promoting skill acquisition through social interaction and psychological support. Patients benefit by collaborating on exercises, sharing challenges, and engaging in motivational or competitive activities [86, 87]. However, group therapy is often limited to seated exercises, as therapists cannot provide individualized physical assistance simultaneously to multiple patients for functional movements such as walking or StS transitions. While seated exercises may be suitable for upper-limb rehabilitation, clinical guidelines emphasize the need for high-intensity gait training for conditions such as stroke and incomplete SCI to improve locomotor function [12].
RHHI could enable multiple patients to safely perform high-intensity functional exercises while engaging with therapists and peers, helping them achieve the necessary training dosage to facilitate neuroplasticity and recovery. However, to the best of our knowledge, no study has virtually connected multiple patients and/or therapists using robotic devices. Aprile et al. [88] explored the feasibility of a robot-assisted rehabilitation area where a single physiotherapist could supervise up to four patients simultaneously. However, the robotic devices were not virtually connected, limiting interactive engagement. Therapists primarily assumed a supervisory role rather than actively interacting with patients, which reduced their ability to assess performance, provide real-time feedback, and adapt interventions to individual needs.
Another potential group therapy application is music-based rehabilitation, where therapists and patients play instruments together. Michalko et al. [89] proposed using RHHI to mediate interactions between expert and novice musicians, providing real-time feedback for posture and finger positioning. This concept could be extended to NR, engaging multiple patients in complex coordinated functional tasks that enhance synchronization and collaboration. The benefits of music therapy go beyond motor synchronization—music itself can enhance emotional engagement, stimulate cognitive processes, and promote neuroplasticity, making it a powerful and motivating tool in rehabilitation contexts [90, 91].
RHHI has the potential to optimize therapist time while actively engaging multiple patients in interactive rehabilitation. Therapists can simultaneously monitor and adjust interventions, providing personalized guidance through robotic systems [92, 93]. By adapting exercises to individual needs within a group setting, RHHI can enhance motivation, ensure appropriate challenge levels, and improve functional outcomes [94]. Further discussion of research directions for this framework is provided in Sec. VII.

### v2

@@ IV RHHI Potential in NR
RHHI combines the advantages of robotics (Sec. IV-A) while incorporating the therapist’s expertise and the human adaptability into the control (Sec. IV-B). This section summarizes both contributions (Tab. I).
@@ IV-A Added Value of Robotics
Robotic devices complement therapists by partially offloading the patient’s weight and accurately transmitting the therapist’s intended movements to the patient [gassert2018rehabilitation]. This feature is particularly valuable during prolonged and intensive rehabilitation sessions that require continuous physical effort. It is especially beneficial for patients with severe motor impairments, as therapists must maintain patient posture, joint alignment, and smooth movements throughout an exercise. Moreover, the high repeatability of robotic devices can relieve therapist effort by replicating movements from previous demonstrations.
@@
In addition, robotic platforms enrich rehabilitation data by capturing simultaneous kinematic and dynamic information from both the patient and the therapist [riener2006human]. Kinematic data provide movement trajectories, joint angles, and velocities, revealing patterns and achievable ranges of motion. Dynamic data, such as interaction forces, account for the level of assistance or resistance provided and patients’ kinetic responses, giving deeper insight into the quality of the interaction.
@@
RHHI holds great potential beyond clinical settings. In recent years, the availability of commercially-available assistive devices intended for domestic use has been increasing [sunny2023assistive]. Applying RHHI to these devices could enable continued support for exercises outside of the clinic. For individuals requiring continuous care, RHHI enables remote assistance and rehabilitation, allowing caregivers and therapists to monitor activities, provide real-time support, and potentially ensure a more effective continuum of care, even in situations where physical proximity is limited, such as for patients with autism [santos2021design, fachantidis2020tauhe, jouaiti2019robot].
IV-B Added Value of Therapists and Other Humans
In contrast to traditional robotic systems, which typically rely on predefined trajectories and often require time-consuming calibration, RHHI allows exercise customization and adjustment on-the-fly by incorporating therapist expertise, human adaptability, and the ability to observe the patient directly. While this would most obviously be done by therapists, who are trained to support patients, it could (with proper robotic support) be administered by caregivers [Baur2018RobotSupportedTraining] or potentially even less severely impaired patients.
@@ IV-B Added Value of Therapists and Other Humans
Moreover, traditional systems often require settings to be adjusted by non-technical users, like clinicians, who may not fully understand their impact, potentially limiting the effectiveness and usability of these systems. Whereas in RHHI, the therapist’s intervention dynamically shapes the robotic response, allowing for real-time adaptation based on patient feedback and performance, enhancing intuition in robotic interventions [hasson2023neurorehabilitation, amato2024unidirectional]. During each session, therapists can flexibly select and modify exercises or movements based on immediate and precise patient assessments without having to set predefined trajectories.
@@
Therapists can also modulate assistance levels throughout rehabilitation. They may initially offer substantial support, provide transparent supervision during assessments, or introduce resistance to encourage active engagement. Although quantifying assistance or resistance can be difficult, therapists can intuitively assess these needs through direct interaction and visual observation, as in conventional therapy [baur2019beam, luciani2024therapists] and caregivers or even patients themselves can potentially perform the adaptation themselves if a therapist is not available [Gorsic2017ComparisonExercises].
@@ V RHHI Applications in NR
The potential applications of RHHI in NR are diverse and far-reaching. This section begins by presenting a potential implementation of RHHI for impairment assessment (Sec. V-A), followed by three different training examples (Sec. V-B, V-C, V-D). Finally, we explore the use of RHHI in group therapy (Sec. V-E) and remote therapy (Sec. V-F). All approaches are supported by pilot studies; however, more statistically powered investigations are needed to assess their actual effectiveness.
@@ V-A Motor Assessment
Assessing patient impairment levels is crucial in rehabilitation [riener2006human, longatelli2023instrumented]. A thorough evaluation must capture four domains: pathology, impairment, participation, and contextual factors [wade2001research]. When such information is recorded in a standardized, well-structured format, this enriches insight into the rehabilitation process, streamlines data sharing and reuse, and ultimately delivers more cost- and time-efficient care [berenspohler2021common]. Assessing physiological changes (Box V-A) in the brain or spinal cord to understand motor recovery is often not clinically feasible. Therefore, proxies such as muscle weakness, spasticity, and motor coordination are often evaluated and used as markers for recovery. Functional task scales, such as the 9-hole peg test and Box & Blocks, offer valuable information about a person’s ability to perform daily activities [stucki2007international]. Because these metrics can change over time, continuous monitoring is crucial to adapt therapy to patient needs.
@@
Combining the quantitative precision of robotic devices with the personalized insights of manual assessments through RHHI enables a comprehensive and adaptive evaluation approach. An innovative RHHI example is the “Beam-Me-In” strategy presented by Baur et al. [baur2019beam]. Here, a therapist and patient are connected through exoskeleton robots, enabling therapists to directly perceive patient limitations (e.g., muscle tone, range of motion). This embodied experience provides therapists with a more accurate understanding of patient impairments compared to traditional observation alone. RHHI simultaneously quantifies joint kinematics and interaction forces, enhancing the precision of impairment assessments and treatment adjustments. At the same time, the therapist can design assessment movements in real time, informed by clinically relevant parameters—such as perceived effort, fatigue, or pain—that are difficult to quantify objectively. Box. V-A visualize how physiological signals can complete the therapist assessment and allow a better understanding of the patient’s impairment during HRRI.
@@
Box 1: Physiological Signal and RHHI Electromyography (EMG) quantifies volitional input during RHHI and reveals how muscles co-contract to stabilize joints and adapt to partner movements and task uncertainty [short2025effects, Takagi2018HapticInteraction, borner2023physically]. In teleimpedance control [peternel2022after], EMG provides real-time muscle activity data, enabling robots to dynamically adjust impedance. Integrating teleimpedance into RHHI could allow bidirectional control of joint posture and user impedance. Combined with Functional Electrical Stimulation (FES), therapist EMG signals could trigger patient stimulation to promote synchronized movement and motor relearning [dalla2024hybrid]. Electroencephalography (EEG) can be used to assess and quantify Interpersonal Synchronization, and simultaneous EEG measurements, known as EEG hyperscanning, have been used to provide insights into neural synchronization between individuals engaged in shared activities [short2021eeg]. Research has shown that inter-brain synchronization, particularly in regions associated with the mirror neuron system [van2009understanding], increases during coordinated activities [dumas2010inter, yun2012interpersonal]. Autonomic Nervous System response, such as respiration rate and skin conductance, offers alternative markers for interpersonal engagement [darzi2021automated, pecoraro2022psychophysiological, tamantini2024fuzzy]. In RHHI, these techniques can provide indicators of motivation and engagement and could potentially be used to, e.g., dynamically adapt the difficulty of a competitive or collaborative exercise for two users with different impairment levels [darzi2021automated].
@@ V-B Upper-Limb Training
Several studies have investigated RHHI during upper-limb training, showing that healthy individuals track movement trajectories more accurately when haptically connected with a partner compared to training alone. These improvements have been linked to estimation of the partner’s movement strategy to refine their own movement control [takagi2017physically, Takagi2018HapticInteraction, Takagi2019IndividualsGoal].
Despite promising findings, few existing RHHI studies have evaluated these approaches in individuals with neurological impairments. Some studies have involved patient populations performing social rehabilitation games [baur2023competitive, Gorsic2017AGame], however, the infrastructure used in these studies did not allow users to physically interact. Both studies emphasized social interaction to boost engagement during training, highlighting the importance of the interaction scenario in future RHHI work. The only notable exception is the study by Waters et al. [waters2024theradyad]. This work presented preliminary results on three chronic stroke participants paired with healthy partners during a trajectory tracking task, showing improved individual motor learning as a result of haptically interacting with an unimpaired user and thus validating the feasibility of such an approach for upper-limb training. Future work could extend these RHHI findings during trajectory tracking tasks to larger cohorts in longitudinal studies to determine the effects of such training on sustained functional benefits.
One promising research direction has been suggested by early studies on human–human interaction during handwriting [buscaglione2024human, way2024design]. To date, RHHI research has primarily focused on arm-level tasks, with limited attention to hand-specific activities requiring dexterous manipulation. Although various hand rehabilitation robots have been developed [ratz2024enhancing, ranzani2023design, kottink2024therapy], none currently support direct haptic interaction between users. Enabling such interaction could significantly expand the range of functional and dexterity-based tasks that patients can perform in collaboration with therapists.
@@ V-C Sit-to-Stand Training
The ability to transition from sitting to standing (StS) is fundamental for independent living. It is one of the most common functional tasks and correlates with clinical balance scores in stroke patients [cheng1998sit]. Post-stroke individuals often struggle with StS due to deficits in strength, motor control, and proprioception, leading to asymmetric force distribution, slower movements, and greater pressure displacement [boukadida2015determinants].
Vianello et al. [vianello2021human] introduced a dual-exoskeleton system that allows a therapist and patient to interact physically across multiple joints and contact points. This setup enables the therapist to guide the patient while incorporating programmable support features, such as balance assistance and safety measures essential for StS tasks. The system was tested with healthy individuals performing synchronized StS movements using shared visual feedback. To reduce fatigue during prolonged sessions, the exoskeletons offload part of the patient’s body weight, allowing them to focus on coordination rather than physical exertion. The study also explored two interaction configurations (Sec. III-A): (1) joint-level coupling, where a spring-damper system links the joint configurations of both users, and (2) task-space coupling, where a virtual connection between their centers of mass allows for flexible posture while preserving overall coordination.
By combining robotic assistance with therapist guidance, RHHI could provide a more adaptive and effective approach to StS training, and could be beneficially extended to standing balance training using a similar setup.
@@ V-D Gait Training
Conventional gait therapy typically involves therapists manually guiding patients to improve balance and gait patterns, often using partial body-weight support. During treadmill-based training, the therapist usually sits in front of the patient and manually assists leg movements. In overground training, body-weight support is also commonly employed, and multiple therapists may be required to support and guide both legs [hesse2000mechanized]. Conventional gait therapy may include tasks such as stair climbing and obstacle avoidance; however, these require specialized body-weight support systems to safely assist patients during more complex maneuvers: equipment that is not available in all clinical settings. Robot-Assisted Gait Training (RAGT) supports patients during high-intensity training [gassert2018rehabilitation, ditommaso2023biomechanics], but it typically facilitates only a limited range of exercises—most often treadmill walking, and less frequently, overground walking.
RHHI offers adaptive, interactive exercises tailored to individual patient needs and the possibility to extend RAGT to a wider range of exercises without prior trajectory definition. Building on this concept, a recent framework virtually physically coupled two healthy participants wearing lower-limb exoskeletons and evaluated how different interaction medium affected inter-user coordination during treadmill walking [kuccuktabak2023virtual]. The same framework was subsequently applied in a physical therapy setting with post-stroke individuals, where the therapist and patient were bidirectionally connected at the hips and knees [kucuktabak2025therapist]. This system allows both sides of the exoskeleton to be actively engaged, enabling natural coordination and force exchange between the therapist and patient during gait training. Moreover, the virtual coupling parameters can be modified to accommodate different therapeutic goals, movement asymmetries, and patient abilities, offering flexibility across training scenarios. When compared with conventional manual therapy, this approach produced superior outcomes, including greater joint range of motion, longer steps, and higher muscle activation, alongside a high level of patient engagement and motivation.
@@
While these results are promising, one limitation of this approach is the need for multiple exoskeletons, which may restrict its practicality and scalability in clinical settings. To address such constraints, alternative designs have explored simplified or asymmetric coupling methods that reduce hardware complexity while maintaining interactive guidance. Amato et al. [amato2024unidirectional] proposed and validated a system in which a therapist wearing an Inertial Measurement Unit (IMU)-based suit provides unidirectional haptic feedback to a patient in an exoskeleton during treadmill walking and obstacle avoidance. Koh et al. [koh2021exploiting] implemented a cross-limb interaction by virtually connecting a robotic manipulandum to a robotic arm physically attached to the paretic leg of post-stroke patients. This setup enables bidirectional interaction, allowing therapists to actively influence the patient’s leg movements and demonstrating improved kinematics during training.
@@
Both approaches [amato2024unidirectional, koh2021exploiting] support collaborative interaction and assist patients in forming internal movement models through therapist-guided feedback. While only Amato et al. demonstrated obstacle avoidance, both systems have the potential to support a wider range of exercises. The main limitation lies in hardware constraints; for example, Koh et al. use a grounded manipulator to assist the paretic leg, which may restrict its use in overground activities such as walking, stairs, or obstacles. Future developments could explore mobile manipulators to extend support beyond treadmill-based training (see Sec. VI for future technical challenges). A key distinction between the two examples proposed is that the method by Amato et al. uses a unidirectional connection—only the patient perceives haptic feedback—which lowers cost and setup time [amato2024unidirectional] (see Sec.VII). Another difference is that the approach of Amato et al. supports interaction with both patient limbs, while Koh et al. focuses on a single limb. Koh et al. tested their system with hemiparetic stroke patients, targeting the impaired limb, but adapting the method for other populations, such as those with SCI, would require further development [koh2021exploiting].
@@ V-E Group Therapy
Group therapy is widely used in rehabilitation, promoting social interaction and psychological support. Patients benefit by collaborating on exercises, sharing challenges, and engaging in motivational or competitive activities [zanca2013group, church2019effectiveness]. However, group therapy is often limited to seated exercises, as therapists cannot provide individualized physical assistance simultaneously to multiple patients for functional movements such as walking or StS transitions. While seated exercises may be suitable for upper-limb rehabilitation, clinical guidelines emphasize the need for high-intensity gait training to improve locomotor function [hornby2020clinical].
RHHI could enable multiple patients to safely perform high-intensity functional exercises while engaging with therapists and peers, helping them achieve the necessary training dosage for recovery. However, to the best of our knowledge, no study has virtually connected multiple patients and/or therapists using robotic devices. Aprile et al. [aprile2019improving] explored the feasibility of a robot-assisted rehabilitation area where a single therapist could supervise up to four patients simultaneously. However, the robotic devices were not virtually connected, limiting interactive engagement. Therapists primarily assumed a supervisory role rather than actively interacting with patients, which reduced their ability to assess performance, provide real-time feedback, and adapt interventions to individual needs.
Another potential group therapy application is music-based rehabilitation, where therapists and patients play instruments together. Michalko et al. [michalko2024enhancing] proposed using RHHI to mediate interactions between expert and novice musicians, providing real-time feedback for posture and finger positioning. This concept could be extended to NR, engaging multiple patients in complex coordinated functional tasks that enhance synchronization and collaboration. The benefits of music therapy go beyond motor synchronization—music itself can enhance emotional engagement, stimulate cognitive processes, and promote neuroplasticity, making it a powerful and motivating tool in rehabilitation contexts [sihvonen2017music, leonardi2018role].
RHHI has the potential to optimize therapist time while actively engaging multiple patients in interactive rehabilitation. Therapists can simultaneously monitor and adjust interventions, providing personalized guidance through robotic systems [miller2022automated, adhikari2023learning]. By adapting exercises to individual needs within a group setting, RHHI can enhance motivation, ensure appropriate challenge levels, and improve functional outcomes [Baur2018TrendsGames]. Further discussion of research directions for this framework is provided in Sec. VII.